If length of E. Coli DNA is 1.36 mm, what is the number of base pairs in E. Coli?

Grade:12

5 Answers

varun
6 years ago
HELLO AKSHAY,
The distance between two consecutive base pairs is 0.34 X 10^-9m
The total length of double helix DNA in an organism can be calculated by multiplying the total number of base pairs with distance between two consecutive base pairs 
Total length of double helix DNA = total number of base pairs X distance between two consecutive bp
 
Let the total number of bp be x
1.36mm = x X 0.34 X 10^-9m
by solving this we get,
= 4 X 10^6 bp
So the number of base pairs in given strand of E coli = 4 X 10^6bp
THANK YOU
Ayce Nur Yucesoy
6 years ago
Distance between two consecutive base pairs is 0.34 × 10^-9 m
To find the number of base pairs, you need to divide the length of the distance.
Darshan
6 years ago
Total length of given DNA = product of total number of basepairs and distance between themThis above formula is applicable for all lengths of DNA provided information is correct
Aashu Mittal
6 years ago
AIPMT 2016 Exam Pattern helps the applicants to understand the outline of the paper well which facilitate to prepare and plan in advance. The pattern specifies the duration of the paper, number of questions, marks allotted per subject, syllabus etc.
Vaibhav Bandu Bargaje
5 years ago
To find the no of Base pairs in E.coli 
 
Given :
Let the length of e.coli DNA is 1.36mm/1.36×10^3m
And distance between two consecutive (bp) 
 is 
0.34 ×^-9m 
 
Let 
Length of DNA =no. of Base pairs ×distance between consecutive (bp ) 
1.36×10^3m =no. Of Base pairs ×0.34×10^-9m
No of Base pairs =1.36×10^3÷0.34×10^-9
By solving it 
No of Base pairs =4×10^6m